Prove that equivalent resistance R of other resistor R1 R2 R3...RN in series is given as

R= R1+R2+R3+...RN


1
Expert's answer
2021-03-10T17:16:10-0500

By the law of the conservation of charge/the continuity of current we find that the current in each element is the same "I_1=I_2=...=I_N=I".

As the tension "U" is equal to the potential difference between the ends of the series of elements, we can decompose it as "U = \phi_{after}-\phi_{before}=(\phi_{after N}-\phi_{after (N-1)})+(\phi_{after (N-1)}-\phi_{after (N_2)})+...+(\phi_{after 1}-\phi_{before 1})" where "\phi_{before/after \text{ }i}" means the potential right before/after the i-th element. Thus, we deduce that

"U= \sum U_i", i.e. the total tension is distributed between the elements.

Now by applying the Ohm's law to each element we have "U_i = I_i R_i", "U=IR_{equivalent}". Applying the two previous results we find "R_{equivalent} = \sum U_i/I = \sum R_i"